OpenStudy (briellshaw):

If I'm 55 and I have 44 Friends Can I Split all my friends into groups for bingo night how many groups could I Have?????

OpenStudy (anonymous):

I don't think your age has anything to do with it lol.

syed98 (syedmohammed98):

umm age has nothing to do with that question...

OpenStudy (igreen):

Unless you're leaving out information.

OpenStudy (igreen):

Any more information given? @BriellShaw

OpenStudy (anonymous):

It's probably just a self made question.

OpenStudy (briellshaw):

Oh my god it was just a random questin the answer is 44/4=11 11 friends in 4 groups it's easy duh

OpenStudy (anonymous):

:)

OpenStudy (briellshaw):

i knew the answer the whole time

OpenStudy (anonymous):

So why you post it.

OpenStudy (briellshaw):

just to do it

syed98 (syedmohammed98):

so u wanted to test Open Study?

OpenStudy (anonymous):

Test people to see how many would get distracted by this spam question.

syed98 (syedmohammed98):

lol

OpenStudy (igreen):

@BriellShaw That's not entirely correct. That is one answer. 44 / 2 = 22 22 groups of 2 people 44 / 4 = 11 11 groups of 4 people And others..

OpenStudy (igreen):

There is more than 1 answer to the question :)

OpenStudy (briellshaw):

STOP CHANGING MY ANSWER SMART ALEK

syed98 (syedmohammed98):

Please post question with full info :). We would be really grateful to you.

OpenStudy (anonymous):

It says in your profile that you're 13 years old.

OpenStudy (anonymous):

So if you're 55 and you're 13 at the same time, what age are you really?

syed98 (syedmohammed98):

take the average i guess.... (13+55)/2

OpenStudy (anonymous):

That's old.

syed98 (syedmohammed98):

yes it is .. so she is 34 ...

OpenStudy (anonymous):

Aha, I knew it was somewhere between 30 and 40.

syed98 (syedmohammed98):

lol :)

OpenStudy (anonymous):

I think it's a transvestite lol.

syed98 (syedmohammed98):

:O....lol

OpenStudy (anonymous):

Is that you in the pic?

syed98 (syedmohammed98):

yupp

OpenStudy (anonymous):

You are hot.

syed98 (syedmohammed98):

thnx :)